OpenStudy (anonymous):

Find y' when y is given by x^11 Is this the correct answer? y' = 10 x^11

OpenStudy (loser66):

nope

OpenStudy (agent0smith):

Multiply by the power FIRST, THEN reduce the power by one you kinda did it the other way around.

OpenStudy (agent0smith):

Eg y' for y = x^8 is y' = 8x^7

OpenStudy (anonymous):

y1*11 11-y1 y' = 11 x10 is that right?

OpenStudy (agent0smith):

if you meant... y' = 11 x^10

OpenStudy (anonymous):

yeah, thanks man.